Solve, finding all solutions in [0,271). 25. 4= 4 sin’ x+3 26. cos” x = 2 cos x 27. 1–3 tan” x = 0

I know the limit is only question per post, so if you could please just help me out with number 25! I havent done this in awhile and I need my memory refreshed with how to do these problems. Thank you!
Answer #1

(25 Given trigonometric equation is 4-4 sins +3 - 1= 4 sinoc & sint oc = & Sintxe = 6 Siña = (sin II) 7 Sim x = Sint Am The
Answer #2

SOLUTION :


25.


4 = 4 sin^2 (x) + 3 

=> 4 - 3 = 4 sin^2 (x)

=> sin^2 (x) = 1/4 

=> sin (x) = +/- 1/2 


So, in the interval [0, 2pi) :


When sin (x) = + 1/2 : 

 

Positive value of sin(x) occurs in 1st and 2nd quadrants.

x = pi/6 , (pi - pi/6)

=> x = pi/6, 5pi/6


When sin (x) = - 1/2 :


Negative value of sin(x) occurs in 3rd and 4th quadrants.

x = (pi + pi/6), (2pi - pi/6)

=> x = 7pi/6, 11pi/6  


So, when sin (x) = +1/2 , - 1/2 

x = pi/6, 5pi/6, 7pi/6, 11pi/6 (ANSWER).

answered by: Tulsiram Garg